What is the mass of sodium iodide in 50.0 mL of 2.63  10–2 M NaI(aq)?


1
Expert's answer
2021-02-09T03:36:10-0500

"No. Moles (mol) = Molarity (M) x Volume (L)"

Convert 50mL into liters

"1L =1000mL\newline ??=50mL\newline =0.05L"

Therefore,

"No\ of\ moles=2.63*10^{-2}*0.05 \newline No\ of\ moles=0.001315mols"


Calculate the mass of sodium iodide

Mass is calculated using the following formula,

"Mass=no\ of\ moles*molar\ mass\newline mass=0.001315*149.89 g/mol
\newline =0.2g"